Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2024 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 31     Entry time: 23 Feb 2024, 23:03
Author: Tokushi Sato 
Group: SPB 
Subject: Shift summary 

X-ray delivery:

stable and flat over 190 pulses per train. 1.1 MHz with ~200 uJ

Optical laser delivery:

NA

Achievements/Observations:

Beam optimization and characterization
Sample injection and data collection for ~10 hours

reasonable hit rates observed but highly fluctuating, making alignment optimisation slow

Issues:
DAQ - restarts needed throughout the shift
Sample injection - copper samples dont give high hit rate and increaseing sample concentration led to clogging
Focus drift - focus was tuned at around 3:30 pm ( from 4 um back to 1.5 um)
-again at 19:30 when we checked again focus was >2microns FWHM returned to 1.5

ELOG V3.1.4-7c3fd00